Calculate the volume of oxygen required to burn 40ml of octane (C8H18) completely. The combustion of octane is ? (2C8H18+25O2 ➡️ 16Co2+18H2O
1
Expert's answer
2020-06-12T07:28:45-0400

2C8H18 + 25O2 = 16CO2 + 18H2O

Octane's density is 0.703 g cm−3

40 mL of it weights:

40 x 0.703 = 28.12 g

M (octane) = 114.23 g/mol

n = 28.12 / 114.23 = 0.246 mol

The mole ratio of octane and oxygen gives us: 25/2 x 0.246 = 3.077 mol of O2

At STP the molar volume of a gas (Standard Temperature and Pressure) is equal to 22.4 L 

3.077 x 22.4 = 68.93 L of O2